8) Pat is mixing two sulfuric acid solutions, one 40% and one 10%. To get 15 liters of a solution that is 30%
sulfuric acid, how much of each must she use?

Answer by Alan3354(69443)

Pat is mixing two sulfuric acid solutions, one 40% and one 10%. To get 15 liters of a solution that is 30%
sulfuric acid, how much of each must she use?
----
f = amount of 40%
t = amount of 10%
---
f + t = 15 --- total solution
40f + 10t = 15*30 --- total acid